OpenStudy (anonymous):

I'm having a bit of a conundrum: Two point charges are separated by 5 cm. The attractive force between them is 20 N. Find the force between them when they are separated by 10 cm.

OpenStudy (anonymous):

Coulomb's law states that \[F= \frac{ C_1*C_2 }{ (D)^2 }\]

OpenStudy (anonymous):

The solution turns to be that the Coulomb force decreases inversely to the square of the distance. Since the distance increased twice, the force will decrease four times, so 20N/4=5N.        

OpenStudy (anonymous):

I need some explaining of that matter that the "the distance increased twice, the force will decrease four times,"

OpenStudy (anonymous):

Whoop, figured out my own question! I took the 10 and squared it getting 100. I suppose i could cancel out the cm and get 5N.

OpenStudy (calculusfunctions):

The force between two electrical charges is directly proportional to the product of their charges and inversely proportional to the square of the distance between them. We could invoke a simple algebraic approach to the problem.\[F _{e}=\frac{ kq _{1}q _{2} }{ r ^{2} }\]where k is the constant of proportionality. Now, when they're 5 cm apart, the force between them is 20 N. Hence\[20=\frac{ kq _{1}q _{2} }{ 5^{2} }\] Hence multiplying this equation by 25 yields\[500=kq _{1}q _{2}\]Thus if r = 10 cm, then\[F _{e}=\frac{ kq _{1}q _{2} }{ 10^{2} }\]However\[kq _{1}q _{2}=500\]from equation 1. Thus\[F _{e}=\frac{ 500 }{ 100 }=5\]∴ when they are 10 cm apart, they have a force of 5 N between them.

OpenStudy (anonymous):

Well.... that makes absolutely more sense than my primitive notion. My brain thanks you kindly!

OpenStudy (calculusfunctions):

You're very welcome! I'm a mathematician and a mathematics teacher. Therefore I invoke math into everything! LOL